the length of a rectangular garden is 1m more than three times the garden's width. if the perimeter of the garden is 34 find its dimensions.
help me thank you your answer would be appreciated :D

Answer by ankor@dixie-net.com(22740)

the length of a rectangular garden is 1m more than three times the garden's width.
L = 3W + 1
:
if the perimeter of the garden is 34 find its dimensions.
2L + 2W = 34
simplify, divide by 2
L + W = 17
Replace L with (3W+1) from the 1st statement equation
(3w+1) + W = 17
4W = 17 - 1
4W = 16
W = 16/4
W = 4 m is the width
:
I'll let you find the Length, check these solutions in the perimeter formula.
: